OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0.

OpenStudy (anonymous):

use the formula

OpenStudy (anonymous):

isn't it x^2+6x-7=0 ? (use formula if it is +7)

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

Well if i know how to use the formula i would have answered the question

OpenStudy (anonymous):

ax^2+bx+c=0 so a=1, b=6 and c=7

OpenStudy (anonymous):

is x^2+6x+7=0 ? if it is! that -3+radical of 2 -3-radical of 2

OpenStudy (anonymous):

use vamgadu's formula

OpenStudy (anonymous):

\[-3\pm \sqrt{2}\]

OpenStudy (anonymous):

Is that correct thats what i got

OpenStudy (anonymous):

yes